OpenStudy (anonymous):

0.3x-0.2y=4 0.5x+0.5y+=85/19 solve by elimination method

OpenStudy (gw2011):

Multiply the first equation by 0.5 and the second equation by 0.2 and you get: 0.15x-0.1y=2=38/19 Use 38/19 because of the second equation 0.1x+0.1y=17/19 Add both equations and you get: 0.25x=55/19 0.25=1/4 (1/4)x=55/19 x=(55)(4)/19 x=220/19 Substitute x=220/19 into one of the equations and you get: 0.5x+0.5y=85/19 0.5(220/19)+0.5y=85/19 110/19+0.5y=85/19 0.5y=(85/19)-(110/19) 0.5y=-25/19 0.5=1/2 (1/2)y=-25/19 y=(-25/19)(2) y=-50/19 Therefore, x=220/19 and y=-50/19

OpenStudy (anonymous):

\[2x-3/8\le-5 or 2x-3/8\ge5\]

OpenStudy (anonymous):

\[\sqrt[4]{64x ^{12}}y ^{14}\] simplify by factoring assume that all expressions under the redical represent nonegative numbers needs an exact anser use radicals if needed